# Seeded Topology Optimization for Commercial Foundry Integrated Photonics

### v1

@@ 3.2 Blurring Filter Study
The blurring filter used in seeded TO directly enables perturbation of the design parameters, making the size and shape of the filter critical in altering the overall device performance. Other filter shapes that convert a binary geometry to gray-scale are compatible with seeded TO, here we demonstrate the performance of the box filter with a user set X and Y size.
@@
The size of the blurring filter significantly impacts the effectiveness of the seeded TO optimization. A small filter has limited ability to perturb the design parameters and therefore limits the design search space to near the seed device. If the filter is too large, small but critical features of the device are obscured and not necessarily recovered in subsequent optimization. An ideal filter size allows for the maximum design region perturbation without eliminating functional features. The size and shape of the optimal filter may vary depending on device and material platform and will require tuning, however a practical starting size is one quarter of the single-mode waveguide width for the material platform being used. Naively, it may appear that the filter size should be on the order of the minimum feature size from the foundry, but in reality, the dimensions are significantly more dependent on the material platform being optimized. For example, silicon nitride designs would require a larger filter size then the devices in this work, but silicon devices from two foundries with different DRC constraints would have similar optimal filter dimensions.
@@
To investigate the effect of filter size on device performance, a blurring filter study was performed on a 2D ultra-compact MMUX designed using traditional TO to create the seed (Fig. 3a). X and Y filter sizes of 62.5, 112.5, and 187.5 nm were selected for the filter study. The MMUX optimized with the $112.5\times 112.5$ nm blurring filter has excellent performance, however the asymmetric $187.5\times 112.5$ nm filter exhibited slightly better performance. This asymmetric filter has increased blur along the X-axis, which likely elongates the hole at the multimode-waveguide interface without eliminating it, enhancing the TE1 conversion efficiency. Other size filters had either reduced performance or DRC errors due to elimination of critical features or minimal design parameter perturbation respectively. Though the choice of filter size is critical to seeded TO performance, there are no other additional hyperparameters introduced to the optimization enabling seamless transition between traditional TO and seeded TO. The original hyperparameters, such as the physical programming bounds and thresholding parameters, have less effect on a seeded TO optimization due to the device topology being binary and require less tuning for an effective optimization.
4 Test Cases
To demonstrate seeded TO we optimize an ultra-compact MMUX [9], a mode converter [26], and a polarization rotator [27], each initially optimized using traditional TO. We also demonstrate the design of a high contrast grating (HCG) reflector initially optimized using a parameter optimization (PO) method [10]. All of these devices are designed for the GlobalFoundries silicon photonics process (GF45CLO).
4.1 Modal Multiplexer
@@
The original design uses both the silicon and polysilicon layers offered by the GF45CLO process to maximize the number of design parameters available in a compact design region. However, after analysis of experimental results, simulations propounded that the polysilicon layer was not required for an optimal MMUX design [8]. Omitting the polysilicon layer from the original MMUX had negligible effect on simulated device performance, therefore that layer was removed before performing seeded TO.
@@
This device requires parallel optimization of both the TE0 and TE1 performance which are defined by the following FOMs:
$FOM_{TE_{0}}=\left(1-\frac{P_{A}}{P_{0}}\right)+\alpha\frac{P_{B}}{P_{0}},\hskip 28.45274ptFOM_{TE_{1}}=\left(1-\frac{P_{B}}{P_{1}}\right)+\alpha\frac{P_{A}}{P_{1}}$ (7)
@@
Both devices were fabricated and measured using the back-to-back test structure shown in Fig. 4a. A standard fiber-array setup was used to perform wavelength scans with a tunable laser on each device using GF45CLO PDK grating couplers for optical input and output. The open-source photonic integrated circuit testing software LabExT was used to automate measurements [31]. Test structures for both MMUX designs are measured across six separate chips from two wafers and the S-parameters are compared to simulation (Fig. 4).
Figure 4: (a) Measurement test structure for the modal multiplexer with labeled ports. (b) Field plot of the first measurement condition with light input through port 1, transmission measured through port 4, and crosstalk measured through port 3. (c) Field plot of the second measurement condition with light input through port 2, transmission measured through port 3, and crosstalk measured through port 4. The simulated and measured S-parameter spectra are plotted for the TO (d) and seeded TO (e) variants.
The S41 transmission encapsulates two passes through the MMUX for the TE0 channel while the S32 transmission encapsulates two passes for the TE1 channel. The S42 and S31 are the same due to reciprocity and capture the crosstalk between the TE0 and TE1 channels.
The traditional TO MMUX required post-TO manual modification to satisfy DRC, which involved expanding the hole near the multimode waveguide to satisfy the minimum area constraint [9]. These manual modifications move the device topology outside of a local optimum, causing a reduction in the TE1 transmission in both simulation and measurement (Fig. 4d). By strict enforcement of DRC constraints through our DRC correction algorithm, seeded TO produces a topology that satisfies DRC with improved performance for both modes compared the the traditional TO design (Fig. 4e).
An asymmetric blurring filter is used in seeded TO for this device with a larger blur along the x-axis. This allows the DRC limited hole to stretch in x while preventing annihilation of the hole via a large y-axis blur, augmenting the mode conversion efficiency for the TE1 case while maintaining DRC compliance. The final seeded TO topology is smoother than the traditional TO device without the periodic ripples that appear throughout the structure. The disparity in TE1 simulated and measured transmission for the TO MMUX may be caused by the lithography process used to fabricate these devices which often smooths structures with sharp curvature and small features resulting in variation of device performance compared to simulation [32, 33, 34]. The seeded TO MMUX has fewer small, jagged features making it more suitable for the lithography process.
4.2 Mode Converter
@@
Like the MMUX, a TE mode converter is critical for any multimode system. The original device was optimized with a TE0 source in the single-mode waveguide and a a TE1 monitor in the output waveguide. The objective function was designed to only maximize TE1 transmission. For the seeded TO device, a crosstalk term was added to the objective function to maximize extinction:
$FOM=\left(1-\frac{P_{1}}{P_{in}}\right)+\alpha\frac{P_{0}}{P_{in}}$ (8)
$\alpha$ is the extinction coefficient (set to 10 for this optimization), $P_{m}$ is the power of the $m^{\text{th}}$ TE mode in the multimode waveguide, and $P_{in}$ is the input power. This ensures the crosstalk of the final device is low while maximizing mode conversion efficiency. Seeded TO was performed using a square averaging filter instead of the asymmetric filter used for the MMUX.
Figure 5: (a) The transmission of the TE0 to TE1 mode converter is measured using test structures with a varying number of cascaded mode converter pairs. The transmission spectrum is measured for each test structure (b) and a linear fit is applied at each wavelength point (c) to find the loss per device (slope). (d) The crosstalk of the mode converter is measured by using a taper and bend to remove any power in higher order modes after the mode converter. The simulated and measured transmission and crosstalk spectra are plotted for the TO (e) and seeded TO (f) variants.
Test structures were designed to measure both the transmission and crosstalk of each mode converter. A standard fiber-array setup was used to perform wavelength scans on each test structure over the same wavelength span as the MMUX. The transmission measurement uses test structures with 0, 1, 2, 4, and 8 mode converter pairs (Fig. 5a). This allows us to apply a linear fit to the transmission vs. number of devices curve at each wavelength point, the slope of the linear fit is the transmission through one device (Fig. 5b,c). For the crosstalk measurement, we measure a mode converter followed by a taper and a bend (Fig. 5d). The taper converts the TE1 mode to a substrate mode that is lost through the bend. Any power in the fundamental mode after the mode converter is sustained through the taper and bend and measured to determine the crosstalk of the mode converter.
Similar to the Modal Multiplexer, the seeded TO mode converter developed much smoother edges than its traditional TO counterpart. The performance of the seeded TO design is also improved due to lower crosstalk and increased transmission in both simulated and measured data (Fig. 5e,f). The measured transmission of the TO device is roughly 0.4 dB lower than simulation. Like the MMUX, the traditional TO mode converter may also have a mismatch in performance due to lithographic smoothing. However, the seeded TO mode converter has slightly higher measured transmission compared to simulated. This indicates that features developed through seeded TO are more suitable to the fabrication process.
4.3 Polarization Rotator
@@
Figure 6: Simulations reveal that TM transmission and TE extinction spectra are uniformly improved for the seeded TO (b) compared to the traditional TO design (a) of a polarization rotator. (c,d) Normalized log-scale field magnitudes show the rotation of the TE mode via the $|E_{y}|$ field (red) to the TM mode via the $|H_{y}|$ field (blue) for both devices. (e) A layer misalignment study was performed by shifting the polysilicon layer from the nominal position in both X and Y. Heat maps are generated for the polarization rotator transmission at 1550 nm normalized to the maximum transmission (marked with an "x") for the TO (f) and the STO (g) polarization rotator.
The objective function used to design the polarization rotator was designed solely to maximize the TE to TM conversion efficiency:
$FOM=1-\frac{P_{TM}}{P_{in}}$ (9)
Where $P_{TM}$ is the power in the TM mode at the output and $P_{in}$ is the input power. Simulation results reveal the seeded TO polarization rotator outperforms the TO polarization rotator in both crosstalk and transmission over C-band (Fig. 6a,b). Test structures for this device have been included in a future multi-process wafer tapeout.
Like the previous devices, seeded TO smoothed many of the ripples present in the traditional TO polarization rotator. The output waveguide is disconnected from the bulk of the TO device whereas the output waveguide is connected in the seeded TO variant. A shape optimization design methodology would not allow for disconnected features to merge, changing the device topology [12, 13]. The improvement made by seeded TO can be seen in the field plots where the magnitude of the $E_{y}$ field is significantly reduced in the output waveguide of the seeded TO variant compared to the original TO device (Fig. 6c,d).
To investigate the robustness of seeded TO, we performed a layer misalignment study on the polarization rotator. The polysilicon layer was shifted in both X and Y with respect to the nominal position. Both devices are far more sensitive to layer misalignment in the Y-direction due to the small device width. The peak transmission for the TO polarization rotator is shifted 20 nm in X and 20 nm in Y from the nominal position whereas the peak transmission for the seeded TO device is shifted -10 nm in X and 10 nm in Y. This indicates that the seeded TO design is more robust to layer misalignment and a stronger local optimum. The smoothing feature of seeded TO appears to increase the robustness of resulting devices without including any robust-based design scheme in the optimization.
4.4 High Contrast Grating Reflector
[…]
Figure 7: (a) Resonator structure measured to determine the reflectance of the high contrast grating (HCG). (b,c) Measured pass-port transmission spectrum of the resonator with marked FSR ($\Delta\lambda_{FSR}$), drop-port FWHM ($\Delta\lambda_{FWHM,d}$), and pass-port minimum transmission ($T_{min,p}$). The simulated and measured reflectance spectra are plotted for the parameter optimized (PO) (d) and seeded TO (e) variants. The measured reflectances at each resonance peak are binned in 5 nm intervals; the mean, min, and max of each bin are shown. (f) Comparison of PO and seeded TO device geometries.
@@ 4.4 High Contrast Grating Reflector
$FOM=1-\frac{P_{r}}{P_{in}}$ (10)
$P_{r}$ is the reflected power and $P_{in}$ is the input power. To measure this device, a resonator was created using a directional coupler and two HCGs as mirrors (Fig. 7a). This device acts as an add/drop ring resonator due to the counter-propagating light in the resonator. The HCG reflectance ($R$) is related to the drop port FWHM ($\Delta\lambda_{FWHM,d}$), the FSR ($\Delta\lambda_{FSR}$), and the directional coupler through-coupling coefficient ($t$) through the following equation [32, 41]:
@@
The derivation for this can be found in the appendix A. Like the previous devices, the resonator pass-port test structure was implemented using the GF45CLO PDK grating couplers for I/O to be measured using a standard fiber array testing setup. The LUNA OVA 5100 was used for this measurement to ensure sufficient wavelength resolution (1.2 pm for the LUNA OVA compared to 10 pm for the tunable laser sweep). Resonant peaks in the transmission spectrum are identified using a peak-finding algorithm and used to calculate the FSR (Fig. 7b). Each peak is isolated and the FWHM is algorithmically calculated to determine the reflectance at each resonance peak (Fig. 7c). Test structures are measured of 5 separate chips from 2 wafer samples, the calculated reflectances are binned in 5 nm intervals, and the mean, min, and max of each bin are calculated and plotted along with the simulated reflectance (Fig. 7d,e).
@@
Seeded TO had limited visible effect on the topology of the HCG, only modifying the hole at the waveguide-grating interface (Fig. 7f). This feature is critical to achieving maximum performance through shaping of the light entering the grating, but is limited by the minimum enclosed area constraint. The seeded TO modifications are demonstrated to improve performance in both simulation and measurement (Fig. 7d,e). This validates the effect that the minimal modifications made by seeded TO have on the performance of fabricated devices. The changes made by seeded TO around the waveguide-grating interface are small enough that a designer may assume they will not be resolved in fabrication. However, the measured improvement of the fabricated device demonstrates small changes like this are critical to device performance.
5 Conclusion

### v5

@@ III.2 Blurring Filter Study
The blurring filter used in seeded TO directly enables perturbation of the design parameters, making the size and shape of the filter critical in altering the overall device performance. Other filter shapes that convert a binary geometry to gray-scale are compatible with seeded TO, here we demonstrate the performance of the box filter with a user-set X and Y size.
@@
The size of the blurring filter significantly impacts the effectiveness of the seeded TO optimization. A small filter has limited ability to perturb the design parameters and therefore limits the design search space to near the seed device. If the filter is too large, small but critical features of the device are obscured and not necessarily recovered in subsequent optimization. An ideal filter size allows for the maximum design region perturbation without eliminating functional features. The size and shape of the optimal filter may vary depending on device and material platform, and will require tuning; however, a practical starting size is one quarter of the single-mode waveguide width for the material platform being used. Since there are only two parameters defining the blurring filter (one if a symmetric filter is used), a parameter sweep can be performed to determine the optimum blurring filter size. Naively, it may appear that the filter size should be on the order of the minimum feature size from the foundry, but in reality, the dimensions are significantly more dependent on the material platform being optimized. For example, silicon nitride designs are longer and have larger features than corresponding silicon designs therefore a larger filter size is required to perturb the device topology Rahim et al. (2017); however, silicon devices from two foundries with different DRC constraints would have similar optimal filter dimensions as the feature sizes are generally the same.
@@
IV Test Cases
@@
Each test case has an objective function depending on the waveguide mode overlap:
$a_{m}^{\pm}=c\int_{A}\,\left[\mathbf{E}^{*}(r)\times\mathbf{H}^{\pm}_{m}+\mathbf{E}^{\pm}_{m}(r)\times\mathbf{H}^{*}\right]\cdot\mathbf{\hat{n}}\,dA$ (9)
where $\alpha^{\pm}_{m}$ is the overlap coefficient of the $m^{th}$ mode for forward (++) and backward (-) directions, $\mathbf{E}(r)$ and $\mathbf{H}(r)$ are the Fourier-transformed total fields, $\mathbf{E}^{\pm}_{m}(r)$ and $\mathbf{H}^{\pm}_{m}(r)$ are the mode profiles for the forward and backward propagating modes, and $c$ is the normalization constant Hammond et al. (2021). The normalization constant is chosen such that:
$|\alpha^{\pm}_{m}|^{2}=\frac{P}{P_{in}}$ (10)
where $P$ is the total power propagating in that particular mode, which is normalized to the input power ($P_{in}$), ensuring the maximum value of $|\alpha^{\pm}_{m}|^{2}$ is 1.
On-chip measured test structures were all created using FotonixTM PDK grating couplers for optical input and output. A standard fiber array setup was used to couple to the test structures. Spectrum plots were measured using either the LUNA optical vector analyzer (OVA) 5100 or a Keysight 8164B tunable laser source with a Koheron PD10R photodiode. The open-source photonic integrated circuit testing software LabExT was used to automate measurements Eppenberger et al. (2021). The grating coupler insertion loss was calibrated out by measuring direct grating-to-grating test structures. Multiple chips across 2 wafer samples were measured for each device.
IV.1 Modal Multiplexer
The use of higher-order optical modes in integrated photonics has many applications, including photonic computation, high extinction/low loss switching systems, and high data-rate communications using mode-division multiplexing Mojaver et al. (2024); Truong et al. (2021); Yang et al. (2022). In these systems, a MMUX or mode converter is required to couple between modes. We use seeded TO to improve an ultra-compact ($\upmum ×\times 33 μ\upmum$) MMUX designed with traditional TO (Fig. 4a) Hiesener et al. . As described in sec. 3.1, this device converts the fundamental mode of two separate waveguides to the TE0 and TE1 modes in a multimode waveguide (Fig. 4b,c) Piggott et al. (2020).
@@ IV.1 Modal Multiplexer
The original design uses both the silicon and polysilicon layers offered by the FotonixTM process to maximize the number of design parameters available in a compact design region. However, after analysis of experimental results, simulations propounded that the polysilicon layer was not required for an optimal MMUX design Pesch et al. . Omitting the polysilicon layer from the original MMUX had a negligible effect on simulated device performance, therefore, that layer was removed before performing seeded TO.
@@
This device requires parallel optimization of both the TE0 ($f_{1}(\mathbf{E})$) and TE1 ($f_{2}(\mathbf{E})$ performance which are defined by the following FOMs:
$\begin{matrix}f_{1}(\mathbf{E})=1-|\alpha_{0,A}^{+}|^{2}+b|\alpha_{0,B}^{+}|^{2}\\ \\ f_{2}(\mathbf{E})=1-|\alpha_{0,B}^{+}|^{2}+b|\alpha_{0,A}^{+}|^{2}\end{matrix}$ (11)
$b$ is the extinction coefficient, $\alpha_{0,A}^{+}$ is the forward propagating fundamental mode coefficient of the single-mode waveguide A (port 1/4), and $\alpha_{0,B}^{+}$ is the forward propagating fundamental mode coefficient of the single-mode waveguide B (port 2/3) Hiesener et al. (Fig. 4a). This is designed to both maximize transmission and minimize the extinction ratio (ER).
Since no PDK device separates TE modes, modal multiplexers are typically measured using a back-to-back structure shown in Fig. 4a Strang (2007); Ha and Carstensen (2024). Though the particular mode in the multimode waveguide cannot be determined, the correct modes are assumed to be excited due to mode orthogonality, low crosstalk, and alignment with simulation. Wavelength scans were performed using a Keysight 8164B tunable laser with a Koheron PD10R photodiode. Test structures for both MMUX designs are measured across six separate chips from two wafers and the S-parameters are compared to simulation (Fig. 4).
Figure 4: (a) Back-to-back measurement test structure for the modal multiplexer with labeled ports. (b) Field plot of the first measurement condition with light input through port 1, transmission measured through port 4, and crosstalk measured through port 3. (c) Field plot of the second measurement condition with light input through port 2, transmission measured through port 3, and crosstalk measured through port 4. The simulated and measured S-parameter spectra are plotted for the traditional TO (d) and seeded TO (e) variants. The dark lines depict the mean values; the light band depicts the worst- and best-case performance across all measured chips.
The S41 transmission encapsulates two passes through the MMUX for the TE0 channel, while the S32 transmission encapsulates two passes for the TE1 channel. The S42 and S31 are the same due to reciprocity and capture the crosstalk between the TE0 and TE1 channels.
The traditional TO MMUX required post-TO manual modification to satisfy DRC, which involved expanding the hole near the multimode waveguide to satisfy the minimum area constraint Hiesener et al. . These manual modifications resulted in a significant reduction in the TE1 transmission in both simulation and measurement for the traditional TO MMUX (Fig. 4d). In contrast, seeded TO significantly improved the TE1 transmission (S32) while maintaining the TE0 transmission (S41) (Fig. 4e). We note that the FOMs $f_{1}$ (optimizing TE0 transmission) and $f_{2}$ (optimizing TE1 transmission) were evenly weighed; therefore, the optimizer focused on minimizing the highest FOM in each iteration ($f_{2}$ for most of the seeded TO run). Through strict enforcement of DRC constraints via our DRC correction algorithm, the seeded TO topology satisfies DRC without manual modifications to the geometry after TO.
Figure 5: (a) The transmission of the TE0 to TE1 mode converter is measured using test structures with a varying number of cascaded mode converter pairs. The transmission spectrum is measured for each test structure (b) and a linear fit is applied at each wavelength point (c) to find the loss per device (slope). (d) The crosstalk of the mode converter is measured by using a taper and bend to remove any power in higher-order modes after the mode converter. The simulated and measured transmission and crosstalk spectra are plotted for the traditional TO (e) and seeded TO (f) variants. The dark lines depict the mean values; the light band depicts the worst- and best-case performance across all measured chips.
After experimenting with different blurring filter dimensions, a rectangular 15-pixel $\times$ 9-pixel asymmetric moving average blurring filter is used in seeded TO for this device with a larger blur along the X-axis. This allows the DRC limited hole to stretch in X while preventing annihilation of the hole via a large Y-axis blur, augmenting the mode conversion efficiency for the TE1 case while maintaining DRC compliance. The final seeded TO topology is smoother than the traditional TO device without the periodic ripples that appear throughout the structure. The disparity in TE1 simulated and measured transmission for the traditional TO MMUX may be caused by the lithography process used to fabricate these devices which often smooths structures with sharp curvature and small features resulting in variation of device performance compared to simulation Chrostowski and Hochberg (2015); Wang et al. ; Xing et al. (2022). The seeded TO MMUX has fewer small, jagged features making it more suitable for the lithography process.
IV.2 Mode Converter
@@
Like the MMUX, a TE mode converter is critical for any multimode system. The original device was optimized with a TE0 source in the single-mode waveguide and a TE1 monitor in the output waveguide. The objective function was designed to only maximize TE1 transmission. For the seeded TO device, to experiment with a modified objective function a crosstalk term was added to the objective function that maximizes extinction:
$f_{1}(\mathbf{E})=1-|\alpha_{1}^{+}|^{2}+b|\alpha_{0}^{+}|^{2},$ (12)
$b$ is the extinction coefficient, $\alpha_{1}^{+}$ is the forward propagating mode coefficient for the TE1 mode, and $\alpha_{0}^{+}$ is the forward propagating mode coefficient for the TE0 mode. This ensures the crosstalk of the final device is low while maximizing mode conversion efficiency. Multiple values of the extinction coefficient were tested for this device (including the original case $b=0$). The performance of each device was similar; however, $b=10$ was slightly better in simulation of both transmission and crosstalk, and subsequently the only mode converter selected for fabrication. Seeded TO was performed using a square 11-pixel $\times$ 11-pixel moving average blurring filter; optimization runs with larger and asymmetrical kernels were initially attempted, yielding worse performance improvement.
Test structures were designed to measure both the transmission and crosstalk of each mode converter. Like the MMUX, wavelength scans were performed using a Keysight 8164B tunable laser with a Koheron PD10R photodiode. The transmission measurement uses test structures with 0, 1, 2, 4, and 8 mode converter pairs (Fig. 5a). This allows us to apply a linear fit to the transmission vs. number of devices curve at each wavelength point; the slope of the linear fit is the transmission through one device (Fig. 5b,c). This transmission is the total power through the device, including the TE0 mode; however, the crosstalk measurement reveals that the TE0 transmitted power is minimal compared to the total transmission. For the crosstalk measurement, we measure a mode converter followed by a taper and a bend (Fig. 5d). The taper converts the TE1 mode to a substrate mode that is lost through the bend. Any power in the fundamental mode after the mode converter is sustained through the taper and bend and measured to determine the crosstalk of the mode converter.
Similar to the Modal Multiplexer, the seeded TO mode converter developed much smoother edges than its traditional TO counterpart. The performance of the seeded TO design is also improved due to lower crosstalk and increased transmission in both simulated and measured data (Fig. 5e,f). The measured transmission of the traditional TO device is roughly 0.4 dB lower than simulation. Like the MMUX, the traditional TO mode converter may also have a mismatch in performance due to lithographic smoothing. However, the seeded TO mode converter has slightly higher measured transmission than simulation. This indicates that features developed through seeded TO are more suitable to the fabrication process.
IV.3 Polarization Rotator
@@
Figure 6: Simulations reveal that TM transmission and TE extinction spectra are uniformly improved for the seeded TO (b) compared to the traditional TO design (a) of a polarization rotator. (c,d) Normalized log-scale field magnitudes show the rotation of the TE mode via the $|E_{y}|$ field (red) to the TM mode via the $|H_{y}|$ field (blue) for both devices. (e) A layer misalignment study was performed by shifting the polysilicon layer from the nominal position in both X and Y. Heat maps are generated for the polarization rotator transmission at 1550 nm normalized to the maximum transmission (marked with an "x") for the traditional TO (f) and the seeded TO (g) polarization rotator.
The objective function used to optimize the polarization rotator was designed solely to maximize the TE to TM conversion efficiency:
$f_{1}(\mathbf{E})=1-|\alpha_{1}^{+}|^{2}$ (13)
Figure 7: (a) Resonator structure measured to determine the reflectance of the high contrast grating (HCG). (b,c) Measured pass-port transmission spectrum of the resonator with marked FSR ($\Delta\lambda_{FSR}$), drop-port FWHM ($\Delta\lambda_{FWHM,d}$), and pass-port minimum transmission ($T_{min,p}$). (d) The measured reflectances for both devices from a single chip. The simulated and measured reflectance spectra from all 5 chips are plotted for the parameter optimized (PO) (e) and seeded TO (f) variants. The measured reflectances at each resonance peak are binned in 5 nm intervals; the mean, min, and max of each bin are shown. (g) Comparison of PO and seeded TO device geometries.
where $\alpha_{1}^{+}$ is the first TM mode coefficient propagating forward. The TO polarization rotator is optimized with seeded TO using a square 11-pixel $\times$ 11-pixel moving average blurring filter for both the silicon and polysilicon layers. Simulation results reveal the seeded TO polarization rotator outperforms the traditional TO polarization rotator in both crosstalk and transmission over C-band (Fig. 6a,b). Test structures for this device have been included in a future multi-process wafer tapeout.
Like the previous devices, seeded TO smoothed many of the ripples present in the traditional TO polarization rotator. The output waveguide is disconnected from the bulk of the traditional TO device whereas the output waveguide is connected in the seeded TO variant. A shape optimization design methodology would not allow for disconnected features to merge, changing the device topology Michaels, Wu, and Yablonovitch (2019); Lalau-Keraly et al. (2013). The improvement made by seeded TO can be seen in the field plots where the magnitude of the $E_{y}$ field is significantly reduced in the output waveguide of the seeded TO variant compared to the traditional TO device (Fig. 6c,d).
To investigate the robustness of seeded TO, we performed a layer misalignment study on the polarization rotator (Fig. 6e-g). The polysilicon layer was shifted in both X and Y with respect to the nominal position. Both devices are far more sensitive to layer misalignment in the Y-direction due to the small device width. The peak transmission for the traditional TO polarization rotator is shifted 20 nm in X and 20 nm in Y from the nominal position whereas the peak transmission for the seeded TO device is shifted -10 nm in X and 10 nm in Y. The optimal layer alignment for the seeded TO variant is closer to the nominal position than the traditional TO variant, indicating that the local optimization behavior of seeded TO yields a stronger local optimum than traditional TO.
IV.4 High Contrast Grating Reflector
@@
$f_{1}(\mathbf{E})=1-|\alpha_{0}^{-}|^{2}$ (14)
$\alpha_{0}^{-}$ is the fundamental mode coefficient propagating backwards. To measure this device, a resonator was created using a directional coupler and two HCGs as mirrors (Fig. 7a). This device acts as an add/drop ring resonator due to the counter-propagating light in the resonator. The HCG reflectance ($R$) is related to the drop port FWHM ($\Delta\lambda_{FWHM,d}$), the FSR ($\Delta\lambda_{FSR}$), and the directional coupler through-coupling coefficient ($t$) through the following equation Chrostowski and Hochberg (2015); Van (2016):
@@
The derivation for this can be found in Supp. C. Like the previous devices, the resonator pass-port test structure was implemented using the FotonixTM PDK grating couplers for I/O to be measured using a standard fiber array testing setup. The LUNA OVA 5100 was used for this measurement to ensure sufficient wavelength resolution (1.2 pm for the LUNA OVA compared to 10 pm for the tunable laser sweep). Resonant peaks in the transmission spectrum are identified using a peak-finding algorithm and used to calculate the FSR (Fig. 7b). Each peak is isolated and the FWHM is algorithmically calculated to determine the reflectance at each resonance peak (Fig. 7c). Test structures are measured across 5 separate chips from 2 wafer samples, the calculated reflectances are binned in 5 nm intervals, and the mean, min, and max of each bin are calculated and plotted along with the simulated reflectance (Fig. 7d,e,f).
@@
Seeded TO had limited visible effect on the topology of the HCG, only modifying the hole at the waveguide-grating interface (Fig. 7g). While the seeded TO variant required no manual modification of the topology to conform to DRC, the PO variant required modification of the corners of the small semicircle hole to conform to the minimum curvature radius constraint. This feature is critical to achieving maximum performance through shaping of the light entering the grating, but is limited by the minimum enclosed area constraint. Though there is some variance in the measured response, the seeded TO modifications are demonstrated to improve performance in both simulation and measurement. Figure 7d shows that the calculated reflectance of the seeded TO variant is typically higher than the PO variant at corresponding resonances. The average measured/simulated reflectance over C-band is 0.963/0.969 for the seeded TO HCG compared to 0.960/0.963 for the PO variant. This validates the effect that the minimal modifications made by seeded TO have on the performance of fabricated devices. The changes made by seeded TO around the waveguide-grating interface are small enough that a designer may assume they will not be resolved in fabrication. However, the measured improvement of the fabricated device demonstrates that small changes like this are critical to device performance.
V Conclusion